# Prioritized-MVBA: A New Approach to Design an Optimal Asynchronous Byzantine Agreement Protocol

### v1

@@ 5.2 Efficiency Analysis
Each sub-protocol and step, except for Sequential-ABBA, has a constant running time. The running time of the proposed protocol is dominated by the Sequential-ABBA sub-protocol, which has an expected constant number of asynchronous rounds.
@@
In all sub-protocols and steps, except for pVCBC and propose steps, each party communicates with all other parties. Thus, the message complexity is $O(n^{2})$. The expected message complexity of the Sequential-ABBA protocol is also $O(n^{2})$.
@@
The communication complexity of each sub-protocol and step is $O(n^{2}(l+\lambda))$, where $l$ is the bit length of input values and $\lambda$ is the bit length of the security parameter. The expected communication complexity of the Sequential-ABBA protocol is also $O(n^{2}(l+\lambda))$.
In conclusion, the Prioritized-MVBA protocol achieves optimal efficiency and correctness by reducing the number of broadcasts and ensuring all security properties of a multi-valued Byzantine agreement protocol. The detailed analysis confirms its resilience, integrity, liveness, and external validity, making it suitable for robust decentralized infrastructures.
@@ 5.3 Case Study
The key contribution of this paper is the reduction in the number of requesting parties and the subsequent utilization of fewer broadcasts to eliminate the expensive sub-components of the classic MVBA protocol. The main challenge is to provide sufficient information to the parties to maintain the protocol’s progress while removing the costly elements. It is crucial to ensure that at least one selected party’s proposal reaches at least $12f+$ 1 parties. This is achieved through the pVCBC step and the propose-recommend step. To illustrate the protocol’s effectiveness, we present a case study in two parts. Section 5.3.1 demonstrates how the protocol achieves the desired properties with the minimum number of nodes a system can have. To further clarify that the protocol can maintain these properties with a larger number of nodes, Section 5.3.2 provides a case study using charts for a system with three times the number of faulty nodes.
@@ 5.3.2 Message Distribution for More Than One Faulty Node
For this study, we assume a total of $n=3f+1=10$ parties, with $f=3$ being faulty. This configuration allows us to explore various message distribution patterns. We require at least one party’s proposal to reach $72f+1=$ 7 parties, with a total of $f+1=4$ selected parties. We examine the impact of different numbers of faulty parties and message distribution patterns. In each figure, the top box indicates selected parties and the number of parties that can recommend the proposal to others (e.g., Party $p_{1}$ can receive 3 recommendations for the first selected party ($R_{1}$), 3 for the second selected party ($R_{2}$), and 2 for the third selected party ($R_{3}$), totaling 7). The bottom box counts the number of recommendations received for each selected party (e.g., $p_{1}$ receives a total of 7 $R_{1}$ recommendations).
@@ 5.4 Analysis of Communication Complexity and Resilience
This section evaluates our protocol’s performance by comparing its communication complexity with classic MVBA protocols and assessing its resilience, termination, and safety properties against committee-based protocols.
5.4.1 Comparison of communication complexity with classic protocols.
@@
Our work focuses on the first polynomial time asynchronous byzantine agreement protocol and removes the expensive sub-component from that protcol. We compare our work with the classic MVBA protocol and the traditional MVBA protocols. Table 1 highlights these comparisons. Our techniques differ from the recent improvements in the MVBA protocol. Therefore, we compare our protocol with the classic protocol and the following improvement. Our protocol differs from the Cachin-MVBA [13] protocol in terms of communication complexity. VABA [23] is a view-based protocol where each view is an instance (a complete execution of the protocol) of the protocol. VABA does not guarantee that the parties will reach an agreement at each instance. In our protocol, parties can reach an agreement on a valid proposal in every instance. We also present a comparison between our proposed protocol and the Dumbo-MVBA [29] protocols. The Dumbo-MVBA protocol is intended for large inputs. It uses the erasure code technique to remove the expensive term. The erasure code first disperses the message and then recovers the message at the end of the agreement, which is expensive in terms of latency and computation and is not worth it for small-size inputs. Our protocol achieves the desired communication complexity (removal of the $O(n^{3})$ term) without the utilization of information dispersal techniques (erasure code) and in one instantiation of the protocol. We can use the information dispersal technique like Dumbo-MVBA in our protocol to reduce the communication complexity to $O(ln+\lambda n^{2})$, but our protocol does not depend on the erasure code to remove the $O(n^{3})$ term.
@@
Table 1: Comparison for performance metrics of MVBA protocols
Protocols Comm. (bit) Word Time Msg. ErasureCode. Instance
Cachin [13] $O(ln^{2}+\lambda n^{2}+n^{3})$ $O(n^{3})$ $O(1)$ $O(n^{2})$ No 1
VABA [23] $O(ln^{2}+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ No $\geq 1$
Dumbo-MVBA [29] $O(ln+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ Yes 1
Dumbo-MVBA* [29] $O(ln+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ Yes 1
Our work $O(ln^{2}+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ No 1
@@ 5.4.2 Comparison of Resilience, Termination, and Safety with Committee-Based Protocols
We compare our work with notable committee-based protocols, specifically focusing on resilience, termination, and safety properties. Table 2 highlights these comparisons. COINcidence [19] assumes a trusted setup and violates optimal resilience. It also does not guarantee termination and safety with high probability (whp). Algorand [20] assumes an untrusted setup, with resilience dependent on network conditions, and does not guarantee termination whp. The Dumbo [11] protocol uses a committee-based approach, but its committee-election protocol does not guarantee the selection of an honest party, thus failing to ensure agreement or termination. Our protocol achieves optimal resilience and guarantees both termination and safety, as our committee-election process ensures the selection of at least one honest party. This guarantees that the protocol can make progress and reach agreement despite adversarial conditions.
@@
Protocols n> Termination Safety
COINcidence [19] 4.5f whp whp
Algorand [20] * whp w.p. 1
Dumbo1 [11] 3f whp w.p. 1
Dumbo2 [11] 3f whp w.p. 1
@@ A.2 Threshold Signature Scheme
[…]
@@ B.1 Comparison with Atomic Broadcast Protocol
As discussed earlier, when the inputs of each party are nearly identical, outputting the requests of $n-f$ parties is not a viable solution. This approach results in higher computational effort without increasing the number of accepted transactions. Table 3 provides a comparison of the communication complexity of our protocol with atomic broadcast protocols. Notably, no atomic broadcast protocol can eliminate the multiplication of $O(n^{3})$ terms. Additionally, atomic broadcast protocols require extra rounds of message exchanges. Here, we focus solely on the communication complexity.
@@
HB-BFT/BEAT0 [5] $O(ln^{2}+\lambda n^{3}logn)$
BEAT1/BEAT2 [42] $O(ln^{3}+\lambda n^{3})$
Dumbo1 [11] $O(ln^{2}+\lambda n^{3}logn)$
Dumbo2 [11] $O(ln^{2}+\lambda n^{3}logn)$
Speeding Dumbo [10] $O(ln^{2}+\lambda n^{3}logn)$
@@ B.2 The Challenge of Classical MVBA Designs
Upon receiving verifiable proof from $n-f$ parties, signaling the completion of the VCBC protocol by the threshold number of parties, a party can initiate the ABBA protocol. However, there exists the possibility that other parties have not received sufficient verifiable proof, or that the adversary manipulates the distribution of proofs in a manner that prevents the majority of ABBA instances from receiving adequate proof.
To address this, each party communicates with others by transmitting an $n$-bit array, indicating receipt of verifiable proof from $n-f$ parties. Upon receipt of $n-f$ verifiable proofs, a party generates a permutation of the parties and invokes ABBA instances based on the order of permutation, ensuring that the number of ABBA instances remains constant on average.
@@ Appendix C Deferred protocols
Construction of the pVCBC
  • Upon invocation of the pVCBC protocol, a party creates a message using the $ID$ and the requests. The party then authenticates the message using its private key and multicasts the message (lines 03-06).
  • Upon receiving a signed message ($msg_{signed}$) from a party $p_{j}$, a party checks whether the sender is a selected party and whether the message is authentic (signed by the sender). If the sender is a selected party and the message is authentic, the party adds its signature share ($sign-share$) to the message, resulting in $\rho$. The party then replies with $\rho$ to the sender (lines 13-16).
  • Upon receiving a signature share $\rho_{k}$ from a party $p_{k}$, a selected party checks the validity of $\rho_{k}$ (ensuring it was added using $p_{k}$’s key). If $\rho_{k}$ is valid, the party adds the signature share $\rho_{k}$ to its set $\Sigma$ (lines 09-11).
  • A selected party waits for $\langle n-f\rangle$ valid signature shares. Upon receiving $\langle n-f\rangle$ signature shares, the party combines them to generate a threshold signature $\sigma$ (proof that the party has sent the same request to at least $\langle f+1\rangle$ honest parties) and returns $\sigma$ to the caller (lines 07-08).
@@
  • Upon invocation of the Permutation protocol, a party generates a coin-share $\sigma_{i}$ for the instance and broadcasts $\sigma_{i}$ to every party, then waits for $12f+$ 1 coin-shares (lines 22-24).
@@
  • When a party receives a coin-share from another party $p_{k}$ for the first time, it verifies the coin-share (ensuring it is from $p_{k}$) and accumulates the coin-share in the set $\Sigma$. The party continues to respond to coin-shares until it has received $12f+$ 1 valid shares (lines 27-29).
@@
  • Upon receiving $12f+$ 1 valid coin-shares, a party uses its $CToss$ function and the collected coin-shares to generate a permutation of the $n$ parties (lines 24-25).
Sequential-ABBA
The purpose of the Sequential-ABBA protocol is to run an agreement loop to agree on a selected party’s proposal. We use the asynchronous binary Byzantine agreement protocol biased towards $1$ for each selected party until the parties reach an agreement on one of the party’s proposals.
Construction of the Sequential-ABBA
  • The Sequential-ABBA protocol takes two arguments: (i) $\Sigma$, the permutation of the selected parties, and (ii) $W$, the list of verifiable proofs the party has received (line 1).
  • Upon invocation of the protocol, a party declares two variables: (i) $l$, an index number to access the selected parties one by one from the array $\Sigma$, and (ii) $b$, initially set to zero to indicate that the parties have not reached any agreement (lines 2-3).
  • While loop: Parties are chosen one after another according to the permutation $\Sigma$ of $\{1,\ldots,f+1\}$. Let $l$ denote the index of the party selected in the current loop (the selected party $p_{l}$ is called the candidate). Each party $p_{i}$ follows these steps for the candidate $p_{l}$ (lines 4-22):
    • Broadcasts a $vote$ message to all parties containing $u_{l}=1$ if party $p_{i}$ has received $p_{l}$’s proposal and verifiable proof (including the proposal in the vote), and $u_{l}=0$ otherwise.
    • Waits for $\langle n-f\rangle$ $vote$ messages but does not count votes indicating $u_{l}=1$ unless a valid proposal from $p_{l}$ has been received—either directly or included in the $vote$ message.
    • Runs a binary validated Byzantine agreement biased towards $1$ (see Algorithm 6) to determine whether $p_{l}$ has properly broadcast a valid proposal. Vote $1$ if $p_{i}$ has received a valid proposal from $p_{l}$ and adds the protocol message that completes the verifiable broadcast of $p_{l}$’s proposal to validate this vote. Otherwise, if $p_{i}$ has received $n-f$ $vote$ messages containing $u_{l}=0$, then vote 00; no additional information is needed. If the agreement loop decides $1$, exit the loop.
  • Upon reaching an agreement, a party returns the threshold signature of the agreed proposal.
  • Upon receiving $\langle ID,vote,party,u_{l},m\rangle$, a party checks whether $u_{l}=1$. If it receives $u_{l}=1$, then it assigns $m_{l}=m$.

### v2

@@ 5.2 Efficiency Analysis
Each sub-protocol and step, except for Sequential-ABBA, has a constant running time. The running time of the proposed protocol is dominated by the Sequential-ABBA sub-protocol. The Sequential-ABBA sub-protocol runs the ABBA protocol biased towards $1$ in an expected constant number of times (Algorithm 4 takes a permutation list of the selected parties as an argument and iterates over that list). Since we take a permutation of the selected parties (see Algorithm 2 line 19-24), the probability of choosing a selected party is an expected constant number. The ABBA protocol also has an expected constant number of asynchronous rounds. Therefore, the running time of the protocol is an expected constant number.
@@
In all sub-protocols and steps, except for pVCBC and propose steps, each party communicates with all other parties. Every party transmit $O(1)$ information to all other parties (See line 8 of Algorithm 1, line 6 of Algorithm 3, lines 14 of Algorithm 5, lines 3, 14, and 22 of Algorithm 2, lines 16 and 18 of Algorithm 4. Each of the broadcasts sends $O(1)$ information). Since $n$ parties send $O(1)$ information to the $n$ parties, the message complexity is $O(n^{2})$. The expected message complexity of the Sequential-ABBA protocol is also $O(n^{2})$.
@@
The communication complexity of each sub-protocol and step is $O(n^{2}(l+\lambda))$, where $l$ is the bit length of input values and $\lambda$ is the bit length of the security parameter. To calculate the communication complexity we use the same approach as message complexity. We observe that in no step a party transmit $O(n)$ information. Thus, the communication complexity is same as message complexity only includes the bit length of the input values and the bit length of the security parameters. The expected communication complexity of the Sequential-ABBA protocol is also $O(n^{2}(l+\lambda))$.
Worst-Case Analysis:
Table 2 presents the analysis of worst-case rounds for each base protocol. The results show that the worst-case rounds are consistent in terms of $f$ across all the base protocols.
@@ 5.3 Case Study
The key contribution of this paper is the reduction in the number of requesting parties and the subsequent utilization of fewer broadcasts to eliminate the expensive sub-components of the classic MVBA protocol. The main challenge is to provide sufficient information to the parties to maintain the protocol’s progress while removing the costly elements. It is crucial to ensure that at least one selected party’s proposal reaches at least $2f+1$ parties. This is achieved through the pVCBC step and the propose-recommend step. To illustrate the protocol’s effectiveness, we present a case study in two parts. Section 5.3.1 demonstrates how the protocol achieves the desired properties with the minimum number of nodes a system can have. To further clarify that the protocol can maintain these properties with a larger number of nodes, Section 5.3.2 provides a case study using charts for a system with three times the number of faulty nodes.
@@ 5.3.1 Message Flow in Different Stages
Figure 4: Message flow without the presence of faulty nodes.
@@
Figure 5: Message flow in the presence of faulty nodes.
@@ 5.3.2 Message Distribution for More Than One Faulty Node
For this study, we assume a total of $n=3f+1=10$ parties, with $f=3$ being faulty. This configuration allows us to explore various message distribution patterns. We require at least one party’s proposal to reach $2f+1=7$ parties, with a total of $f+1=4$ selected parties. We examine the impact of different numbers of faulty parties and message distribution patterns. In each figure, the top box indicates selected parties and the number of parties that can recommend the proposal to others (e.g., Party $p_{1}$ can receive 3 recommendations for the first selected party ($R_{1}$), 3 for the second selected party ($R_{2}$), and 2 for the third selected party ($R_{3}$), totaling 7). The bottom box counts the number of recommendations received for each selected party (e.g., $p_{1}$ receives a total of 7 $R_{1}$ recommendations).
@@ 5.4 Analysis of Communication Complexity and Resilience
This section evaluates our protocol’s performance by comparing its communication complexity with classic MVBA protocol and assessing its resilience, termination, and safety properties against committee-based protocols.
5.4.1 Comparison of Asynchronous Rounds in MVBA Protocols
We compare the asynchronous rounds required to reach agreement in our pMVBA protocol with those in classic and traditional MVBA protocols, considering both best-case and worst-case scenarios. To determine the worst-case cost, we first calculate the best-case scenario, identify the concrete rounds, and then estimate the remaining rounds as average ABBA rounds, multiplying by $f+1$. As shown in Table 2, Dumbo-MVBA requires the most rounds in the best-case scenario. In the worst-case, while all protocols require rounds proportional to $f$, VABA struggles to guarantee agreement on a complete output after $f+1$ instances, requiring 13 rounds per instance—twice as many as other protocols. Among existing MVBA protocols, pMVBA offers superior performance in communication complexity and asynchronous rounds. Although Speeding Dumbo achieves lower asynchronous rounds, it is an ACS protocol and cannot avoid the $O(\lambda n^{3}\log n)$ communication complexity.
Table 2: Asynchronous rounds of MVBA protocols
Protocols Rounds (Best case) Rounds (Worst case)
Cachin et al. [SECURE03] $\boldsymbol{12}$ $\geq 5+7(f+1)$
VABA [BYZ17] 1313 $\geq 13(f+1)$
Dumbo-MVBA [BYZ20] 1818 $\geq 11+7(f+1)$
Our work 1313 $\boldsymbol{6+7(f+1)}$
5.4.2 Comparison of communication complexity with MVBA protocols.
@@ 5.4.1 Comparison of Asynchronous Rounds in MVBA Protocols
Our work focuses on the first polynomial-time asynchronous Byzantine agreement protocol and removes its expensive sub-component. We compare our work with the classic and traditional MVBA protocols, as highlighted in Table 3. Our techniques differ from recent improvements in MVBA protocols, so we compare our protocol with the classic protocol and the following improvement. Our protocol differs from the Cachin-MVBA [SECURE03] in communication complexity. VABA [BYZ17] is a view-based protocol where each view is a complete protocol execution. VABA does not guarantee agreement in every instance, whereas our protocol ensures parties reach an agreement on a valid proposal in each instance.
We also present a comparison between our proposed protocol and the Dumbo-MVBA [BYZ20] protocols. The Dumbo-MVBA protocol is intended for large inputs. It uses the erasure code technique to remove the expensive term. The erasure code first disperses the message and then recovers the message at the end of the agreement, which is expensive in terms of latency and computation and is not worth it for small-size inputs. Our protocol achieves the desired communication complexity (removal of the $O(n^{3})$ term) without the utilization of information dispersal techniques (erasure code) and in one instantiation of the protocol. We can use the information dispersal technique like Dumbo-MVBA in our protocol to reduce the communication complexity to $O(ln+\lambda n^{2})$, but our protocol does not depend on the erasure code to remove the $O(n^{3})$ term. We use a committee to remove the $O(n^{3})$ term it takes only one message and the cryptographic computation. The cryptographic computation is similar to other parts of the protocol.
@@
Table 3: Comparison for performance metrics of MVBA protocols (E. Code stands for Erasure Code and Inst stands for Instance)
Protocols Comm. (bit) Word Time Msg. E. Code. Inst
Cachin [SECURE03] $O(ln^{2}+\lambda n^{2}+n^{3})$ $O(n^{3})$ $O(1)$ $O(n^{2})$ No 1
VABA [BYZ17] $O(ln^{2}+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ No $\geq$ 1
Dumbo-MVBA [BYZ20] $\boldsymbol{O(ln+\lambda n^{2})}$ $O(n^{2})$ $O(1)$ $O(n^{2})$ Yes 1
Our work $O(ln^{2}+\lambda n^{2})$ $\boldsymbol{O(n^{2})}$ $\boldsymbol{O(1)}$ $\boldsymbol{O(n^{2})}$ $\boldsymbol{No}$ $\boldsymbol{1}$
@@ 5.4.3 Comparison of Resilience, Termination, and Safety with Committee-Based Protocols
We compare our work with notable committee-based protocols, specifically focusing on resilience, termination, and safety properties. Table 4 highlights these comparisons. COINcidence [BYZ19] assumes a trusted setup and violates optimal resilience. It also does not guarantee termination and safety with high probability (whp). Algorand [BYZ21] assumes an untrusted setup, with resilience dependent on network conditions, and does not guarantee termination whp. The Dumbo [FASTERDUMBO] protocol uses a committee-based approach, but its committee-election protocol does not guarantee the selection of an honest party, thus failing to ensure agreement or termination. Our protocol achieves optimal resilience and guarantees both termination and safety, as our committee-election process ensures the selection of at least one honest party. This guarantees that the protocol can make progress and reach an agreement despite adversarial conditions.
@@
Protocols n>> Termination Safety
COINcidence [BYZ19] 4.5f whp whp
Algorand [BYZ21] * whp w.p. 1
Dumbo1 [FASTERDUMBO] 3f whp w.p. 1
Dumbo2 [FASTERDUMBO] 3f whp w.p. 1
@@
6 Implementation and Evaluations
We implemented and evaluated pMVBA protocol using a self-developed gRPC based protocol simulator. Along the way, we also implemented and made systematic comparisons with several typical MVBAs, including VABA [BYZ17] and CKPS01 [SECURE03].
6.1 Performance of pMVBA Protocol
We extensively evaluated and compared our proposed pMVBA protocol with VABA and Classic MVBA (CKSP01) protocol.
Implementation details.
All protocols are written as a single-process Python 3 program. The p2p channels among nodes are established using gRPC insecure channel. The Python multiprocessing library handles the concurrent tasks. Different from [HoneyBadgerBFT] that optimistically skips the verification for threshold signatures, our tests verify all signatures (simulate time to calculate threshold-signatures), which better reflects the actual performance in the worst cases, e.g., with corruptions.
Test environment.
The experiments are conducted using gRPC instances with a similar configuration. All tests scale up to $10$ instances. A transaction in our tests is a string of 1 KB, which approximates the size of 4 basic Bitcoin transactions with one input and two outputs. The batch size (also called the system load sometimes) represents the number of transactions proposed by all nodes in a one-shot agreement, and will vary from 1 to $5\times 10^{2}$.
Result highlights.
The key information from our experiments is that the pMVBA significantly reduces the latency of Classic MVBA (CKSP01) with the increase of system sizes and attains a throughput $72\%$ higher than CKSP01, almost regardless of the system size, cf. Table 5;
Table 5: Improvements of latency and throughput
Peak throughput (tx/s) Basic Latency (sec)
Scale CKSP01 pMVBA CKSP01 pMVBA
[…]
Figure 12: Latency (s) vs Number of Nodes
Basic latency.
We set the input batch size as $4$ transactions (i.e., nearly zero) to test the basic latency of pMVBA (and VABA and CKSP01) and show the results in Fig. 12. It is easy to see that the improvement is consistent with the number of nodes. Though the basic latency of VABA, CKSP01, and pMVBA is increasing with larger system scales; the improvement made by pMVBA is always significant.
Throughput at varying scales.
We then evaluate the throughput of pMVBA at different system scales with varying batch sizes. Along the way, we also compare pMVBA to VABA and CKSP01 and visualize the comparison results in Fig. 13.
As illustrated in the first three subgraphs in Fig. 13, pMVBA attains multi-fold improvements relative to CKSP01, disregarding the system scale and batch size. First, pMVBA directly inherits the advantages of using the 1-step VCBC protocol, after that, using fewer VCBC instances leads to fewer cryptographic computations and message delay time, which helps it finally achieve a throughput twice that of CKSP01 in the larger batch sizes and larger number of nodes. Moreover, carefully analyzing the Fig. 13, we notice that pMVBA noticeably outperforms both VABA and CKSP01 with larger batch sizes and with the increasing number of nodes. The performance improvement is due to the smaller number of computations and messages needed for the pMVBA protocol.
The peak throughputs of all MVBA protocols are comprehensively shown in the last subgraph of Fig. 13. Although the throughput of all protocols decreases with scaling up, pMVBA always maintains the highest peak throughput among them.
Figure 13: Throughput of Several MVBA Protocols
Throughput-latency trade-off.
Fig. 14 shows the throughput latency trade-offs in pMVBA, CKSP01, and VABA. All the trends are roughly straight-line, indicating that all of them eventually become network-bound. pMVBA not only achieves higher peak throughput, but also has a latency that is always smaller than that of VABA and CKSP01 at the same throughput. Moreover, the latency gap between pMVBA and VABA is greatly narrowed at some small scales (VABA) and large scales (pMVBA). This evidence indicates that our multi-fold improvements greatly expand the applicability of the asynchronous BFT protocol.
Figure 14: Latency (ts) vs throughput (tx/s) of the MVBA Protocols.
@@ A.2 Threshold Signature Scheme
  • Key generation algorithm: KeySetup($\{0,1\}^{\lambda},n,f+1)\rightarrow\{UPK,PK,SK\}$. Given a security parameter $\lambda$, this algorithm generates a universal public key $UPK$, a vector of public keys $PK:=(pk_{1},pk_{2},\ldots,pk_{n})$, and a vector of secret keys $SK:=(sk_{1},sk_{2},\ldots,sk_{n})$.
@@ B.1 Comparison with Atomic Broadcast Protocol
As discussed earlier, when the inputs of each party are nearly identical, outputting the requests of nfn-f parties is not a viable solution. This approach results in higher computational effort without increasing the number of accepted transactions. Table 6 provides a comparison of the communication complexity of our protocol with atomic broadcast protocols. Notably, no atomic broadcast protocol can eliminate the multiplication of $O(n^{3})$ terms. Additionally, atomic broadcast protocols require extra rounds of message exchanges. Here, we focus solely on the communication complexity.
@@
HB-BFT/BEAT0 [HONEYBADGER01] $O(ln^{2}+\lambda n^{3}logn)$
BEAT1/BEAT2 [DUAN18] $O(ln^{3}+\lambda n^{3})$
Dumbo1 [FASTERDUMBO] $O(ln^{2}+\lambda n^{3}logn)$
Dumbo2 [FASTERDUMBO] $O(ln^{2}+\lambda n^{3}logn)$
Speeding Dumbo [SPEEDINGDUMBO] $O(ln^{2}+\lambda n^{3}logn)$
@@ B.2 The Challenge of Classical MVBA Designs
Upon receiving verifiable proof from nfn-f parties, signaling the completion of the VCBC protocol by the threshold number of parties, a party can initiate the ABBA protocol. However, there exists the possibility that other parties have not received sufficient verifiable proof, or that the adversary manipulates the distribution of proofs in a manner that prevents the majority of ABBA instances from receiving adequate proof.
To address this, each party communicates with others by transmitting an $n$-bit array, indicating receipt of verifiable proof from nfn-f parties. Upon receipt of nfn-f verifiable proofs, a party generates a permutation of the parties and invokes ABBA instances based on the order of permutation, ensuring that the number of ABBA instances remains constant on average.
@@ Appendix C Deferred protocols
  • Upon invocation of the Permutation protocol, a party generates a coin-share $\sigma_{i}$ for the instance and broadcasts $\sigma_{i}$ to every party, then waits for $2f+1$ coin-shares (lines 22-24).
@@
  • When a party receives a coin-share from another party $p_{k}$ for the first time, it verifies the coin-share (ensuring it is from $p_{k}$) and accumulates the coin-share in the set $\Sigma$. The party continues to respond to coin-shares until it has received $2f+1$ valid shares (lines 27-29).
@@
  • Upon receiving $2f+1$ valid coin-shares, a party uses its $CToss$ function and the collected coin-shares to generate a permutation of the $n$ parties (lines 24-25).
Appendix D Agreement protocol
D.1 Asynchronous Binary Byzantine Agreement (ABBA)
The ABBA protocol allows parties to agree on a single bit $b\in\{0,1\}$ [BYZ10, SECURE05, SIG01]. We have adopted the ABBA protocol from [SECURE02], as given in Algorithm 6. The expected running time of the protocol is $O(1)$, and it completes within $O(k)$ rounds with probability $1-2^{-k}$. Since the protocol uses a common coin, the total communication complexity becomes $O(kn^{2})$. For more information on how to realize a common coin from a threshold signature scheme, we refer interested readers to the [HONEYBADGER01].
Construction of the ABBA biased towards 1
We use the ABBA protocol from [SECURE02]. We optimize and changed the protocol for biased towards $1$. The biases towards $1$ property ensures that if at least one party input $1$ in the pre-process step. The pseudocode of the ABBA protocol biased towards 1 is given in Algorithm 6, and a step-by-step description is provided below:
  • Pre-process step . Generate an $\sigma_{0}$ share on the message and send the pre-process type message to all parties.
  • Collect $2f+1$ proper pre-processing messages. (see (Algorithm 6)).
  • Repeat loop: Repeat the following steps 1-4 for rounds round = 1,2,3,…
    • Pre-Vote step. (see Algorithm 7)
        If round = 1, $b=1$ if there is a pre-processing vote for $1$ (biased towards 1, taking one vote instead of majority) else $b=0$. (see lines 3-4).
        If round >> 11, if there is a threshold signature on main-vote message from round-1 then decide and return. (see lines 18-20)
        Upon receiving main-vote for $0/1$, update $b$ and the justification. (see lines 12-17)
        b=F(ID,r1)b=F(ID,r-1), all the main-vote are abstain and the justification is the threshold signature of the abstain vote. (see lines 6-7)
        Produce signature-share on the message (ID, pre-vote, round, b) and send the message of the form pre-vote,round,b,justification, signature-share). (lines 9-11)
    • Main-vote step. (See Algorithm 8)
        Collect (2f+1) properly justified round pre-vote messages. (lines 14-19)
        If there are (2f+1) pre-votes for 0/1, $v=0/1$ and the justification is the threshold-signature of the the sign-shares on pre-vote messages. (lines 5-7)
        If there are (2f+1) pre-votes for both 0 and $1$, $v=abstain$ and the justification is the two sign-shares from pre-vote 0 and pre-vote 1. (lines 9-10)
        Produce signature-share on the message (ID, pre-vote, round, v) and send the message of the form (main-vote,round,v,justification, signature-share) to all parties (lines 11-13)
    • Check for decision. (See Algorithm 9)
        Collect (2f+1) properly justified main-votes of the round roundround. (line 3)
        If these is no abstain vote, all main-votes for $b\in\{0,1\}$, then decide the value $b$. Produce a threshold signature on the main votes’ sign-shares and send the threshold signatures to all parties and return. (lines 4-7)
        Otherwise, go to Algorithm 9. line (11)
    • Common Coin. (See Algorithm 10)
        Generate a coin-share of the coin (ID, round) and send to all parties a message of the form (coin, round,coin-share). (lines 1-4)
        Collect (2f+1) shares of the coin (ID,round $\sigma_{k}$), and combine these shares to get the value $F(ID,round)\in\{0,1\}$. (lines 5-6)
1
2
3upon ABBA(m) do
   /* Preprocess Step. */
4
5  $\sigma_{0}\leftarrow SigShare_{i}(sk_{i},m_{i})$ see [SECURE02]
6
7  send $(pre$-$process,m_{i},\sigma_{i})$ to all parties
8
9  wait until at least (nf)(n-f) pre-process messages have been received.
10
11  for $round\leftarrow 1,2,3,$... do
12     Prevote Step : Algorithm 7
13
14    Main-vote Step: Algorithm 8
15
16    Check For Decision : Algorithm 9
17
18    Common Coin: Algorithm 10
19
20
Algorithm 6 ABBA biased towards 1: protocol for party $p_{i}$
[…]
2 $justification\leftarrow\perp$
3 if round = 1 then
4   $b\leftarrow 1$ if there is any pre-process message with $m=1$ (biased towards 11), otherwise 0.
5   else
6     $b\leftarrow F(ID,round-1)$
7     $justification\leftarrow threshold$-$signature\langle ID,main$-$vote,round-1,abstain\rangle$
8
9     wait for nfn-f justified main-vote
10
11
12    $m_{i}\leftarrow(ID,pre$-$vote,round,b)$
13     $\sigma\leftarrow SigShare_{i}(sk_{i},m_{i})$
14
15    send $(pre$-$vote,r,b,justification,\sigma)$ to all parties
16
17    upon receiving $\langle main$-$vote,round,v,justification,\sigma\rangle$ for the first time from party $p_{k}$ do
18       if v = 0 then
19         $b\leftarrow 0$
20
21         else if v = 1 then
22           $b\leftarrow 1$
23
24
25          $justification\leftarrow threshold$-$signature\langle ID,pre$-$vote,round-1,b\rangle$
26
27          upon receiving $\langle b,threshold$-$signature\rangle$ for the first time from party $p_{k}$ do
28             send(b, threshold-signature) to all parties
29             return(b, threshold-signature)
30
31
Algorithm 7 ABBA biased towards 1: Pre-vote step
1
2
/* Mainvote Step. */
3
[…]
5 $PV_{0}\leftarrow\{\}$
6 $PV_{1}\leftarrow\{\}$
7 wait until $|\Sigma|$ = 2f+1
8
9if $|PV_{0}|=2f+1$ or $|PV_{1}|=2f+1$ then
10   $v\leftarrow 0/1$
11   $justification\leftarrow CombineShare_{i}(v,{i,\sigma_{i}}_{i\in\Sigma})$
12
13   else
14     $v\leftarrow abstain$
15     $justification\leftarrow(\sigma_{i}\in PV_{0},\sigma_{j}\in PV_{1})$
16
17
18    $m_{i}\leftarrow(ID,main$-$vote,round,v)$
19     $\sigma\leftarrow SigShare_{i}(sk_{i},m_{i})$
20
21    send $(main$-$vote,round,v,justification,\sigma)$ to all parties
22
23    upon receiving $\langle pre$-$vote,round,b,justification,\sigma\rangle$ for the first time from party $p_{k}$ do
24       if b = 0 then
25         $PV_{0}\leftarrow PV_{0}+1$
26
27         else if b = 1 then
28           $PV_{1}\leftarrow PV_{1}+1$
29
30           $\Sigma\leftarrow\Sigma\cup\{\sigma_{k}\}$
31
32
Algorithm 8 ABBA biased towards 1: Main-Vote step
1
2
[…]
4 $isAbstain\leftarrow no$
5 wait until $|\Sigma|$ = 2f+1
6 if isAbstain = no then
7   threshold-signature = $CombineShare_{i}(b,{(i,\sigma_{i})}_{i\in\Sigma})$
8   send(threshold-signature) to all parties
9   return (b, threshold-signature)
10
11   else
12     go to Algorithm 10
13
14
15    upon receiving $\langle main$-$vote,r,v,justification,\sigma\rangle$ for the first time from party $p_{k}$ do
16       if v = abstain then
17         $isAbstain\leftarrow yes$
18
19         $b\leftarrow v$
20         $\Sigma\leftarrow\Sigma\cup\{\sigma_{k}\}$
21
22
Algorithm 9 ABBA biased towards 1: Check for Decision for party $p_{i}$